SX AO Science

00:56
  • Almost finished opening.
  • Finishing up some darks.
02:17
  • Finished seeing limited.
  • Checking pointing.
02:37
  • Preset
  • Same field as las night
  • Measured mag is 13.3
  • Bin 2, 400Hz
  • Gain: 2.55 1.18 0.85
  • Datasets:
  • 024256
  • 025237
03:01
  • Shell riped
  • My Guess, too much TT.
  • The edge rolled and the the shell ripped, instead of the loop opening
03:07
  • Re-sending the preset
  • The TTOffset makes the Camera lens converge much more slowly.
  • Having network slowdown. Not a problem, yet.
  • Lowered TT gain to 2.35
03:12
  • Datasets:
  • 031256
03:20
  • Problems with large offset, the pupils were more than 0.4 pixels off.
  • The offset started before the loop paused again.
  • I had to manually move the stage.
  • We've re-sent the offset.
03:29
  • Dataset:
  • 033013
  • 034044
  • 035041
03:51-ish
  • Had as safe-skip after an offset, for no apparent reason.
04:00
  • Dataset:
  • 040011
04:02
  • An edge started rolling and we weren't able to pause the script before a large offset.
  • The shell ripped
04:09
  • Some large offsets. Value out of range error.
04:19
  • Fighting the offsets, lost 10 min.
04:21
  • Problem with the offsets. They keep saying that we are out of range, when we are not.
  • should have re-sent the preset at 04:09

    Mon Jan 29 04:19:39.681 2018 aos.OffsetXY{FLAO}.started - OffsetXY {FLAO}[477] (ofsx:25.1587[25.1587],ofsy:0.60528)

    Mon Jan 29 04:19:39.721 2018 aos.OffsetXY{FLAO}.failed - OffsetXY {FLAO}[477] - [AOException] RETRY: Offset outside FoV limits. Target xyz is -118.99 -32.71 53.00 (-5002) VALUE_OUT_OF_RANGE_ERROR

04:27
  • Dataset:
  • 042811
  • Stuck in pause
  • Mon Jan 29 04:31:32.676 2018 aos.Resume{FLAO}.started - Resume{FLAO}[327]

    Mon Jan 29 04:31:32.694 2018 aos.Resume{FLAO}.failed - Resume{FLAO}[327] - Illegal command for state LoopClosed

  • The script wasn't resuming (after like a minute), so I manually resumed, after which the loop tried to resume.
04:38
  • re-starting the script, it finished.
  • Datasets:
  • 044222
04:45
  • Top left edge keeps trying to roll (10:00 position)
  • Lowering gains this time
  • 2.00 0.98 0.65
  • Datasets:
  • 045344
  • Now the left edge rolled
  • 050343
  • 051518
05:46
  • Seeing has gotten bad again.
We lost maybe an hour or more due to problems.

-- %USERSIG{GregoryTaylor - 2018-01-29}%
Topic revision: r3 - 29 Jan 2018, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback